A bead of mass M can slide on a thin vertical rod,with sliding friction coefficient between them=z. The rod is translated horizontally with a constant acceleration (a). For what value of (a) will an earth observer see equal horizotal and vertical components of acceleration of the bead?

Grade:

1 Answers

sanki kurli
13 years ago

the bead is sliding  vertically down with friction opposing its direction...

so net force on the bead vertically= M*g - z*N         ....N=normal reaction force=M*a

& net horizontal foce = M*a

the normal foce would be..N=M*a    (as N is in horizontal direction)

according tge cquestion.. the horizontal acce. =vertical acce.

                                      =>(M*g - z*N ) /M = (M*a)/M

                                        =>g - z*a = a

                                         =>a =g/(1+z)